OpenStudy (anonymous):

Use the order of operations to simplify the expression. 16-2 with 3 cubed divided by (2x1)? A.2 B.4 C.8 D.12

OpenStudy (luffingsails):

Not sure what your equation is ibarwick. Can you re-write please?

OpenStudy (anonymous):

the equastion is sixteen minus two to the third power divided by the quantity two times one!

OpenStudy (luffingsails):

\[16-\frac{2^3}{2*1}\] Like this? Trying to be clear because without order of operations I will probably be leading you astray.

OpenStudy (anonymous):

i know but it's multiple chorce (ohh im so confused)!

OpenStudy (luffingsails):

Just trying to get the original equation right, so I can work to solve it. Is that I wrote (the equation) right?

OpenStudy (anonymous):

yes

OpenStudy (luffingsails):

Okay, so let's try and expand it and then simplify. \[16-\frac{2^3}{2*1} = 16 - \frac{2*2*2}{2}\] \[16-\frac{8}{2} = 16-4=12\] Make sense?

OpenStudy (anonymous):

ohh ok